OpenStudy (precal):

Maximizing profit: Suppose you own a tour bus and you book groups of 20 to 70 people for a day tour. The cost per person is $30 minus $0.25 for every ticket sold. If gas and other misc costs are $200, how many tickects should you sell to maximize your profit? Treat the number of tickets as a nonnegative real number.

OpenStudy (precal):

Optimization problem

OpenStudy (rulnick):

c(x)/x = 30 - 0.25x C(x) = c(x) + 200 = 30x - 0.25x^2 + 200 Right? If so, C'(x) = 30 - 0.5x C'(x) = 0 at x = 60. Answer: 60. C''(x) < 0 at x=60 so it is indeed a max.

OpenStudy (precal):

why write c(x) over x?

OpenStudy (rulnick):

Cost per person, right?

OpenStudy (precal):

ok, not sure (you are probably correct)

OpenStudy (precal):

ok if I just check c(x)=[30-.25x](x) 1 person the cost would be 29.75 2 persons the cost should be 59.50 but according to the way we set it up is is 59

OpenStudy (precal):

shouldn't it be c(x)=30x-.25x ?

OpenStudy (precal):

I am trying to understand the problem, process and eventually the solution, any assistant is greatly appreciated

OpenStudy (precal):

Good day AccessDenied ! :)

OpenStudy (rulnick):

I don't think so, not the way you the problem is written. But perhaps I misunderstand it.

OpenStudy (accessdenied):

my understanding is that, you have an income based off of your tickets sold (a revenue), the costs for what you need (cost), and then you just create a profit function based off of it and get that maximum hello. :)

OpenStudy (precal):

well if I did not know any calculus, I would have to say unless I booked 7 people I am operating at a loss

OpenStudy (rulnick):

And, yes, I think the "2 persons" example you gave above it should be 59. Why do you think 59.50?

OpenStudy (precal):

20 people would bring me a profit of 395 70 (my ideal number) would bring me 1882.50 just not sure how to answer problem using calculus (optimization)

OpenStudy (precal):

2 times 30 = 60 but minus .25 per ticket 60-.50=59.50

OpenStudy (rulnick):

I think that maybe ignores the "cost per person." Yours there would be linear cost (cost proportional to number of people). By the way, a correction: cost, not profit, is max at x=60. Cost is min at x=20!

OpenStudy (accessdenied):

Our domain of x: 20 <= x <= 70 Revenue: R(x) R(x)/x = 30 - 0.25x R(x) = 30x - 0.25x^2 Costs: C(x) C(x) = 200 P(x) = R(x) - C(x) = (30x - 0.25x^2) - 200 P'(x) = 30 - 0.5x Evaluate P(x) for zero of derivative + endpoints That's what I'm thinking

OpenStudy (rulnick):

Everything in the problem is expressed as cost, with income presumably appearing implicitly as negative cost. In this case, there is a min profit (max cost) at x=60, but a min cost at the (lower) boundary on tickets sold (always between 20 and 70). Min cost in this range is at x=20.

OpenStudy (precal):

Yes, I recall with rationals we would take the linear equation and place it over x to spread that 200 per person ie the more people that went on the trip reduce the 200 paid per person C(x) = (30 -.25x -200)/x

OpenStudy (rulnick):

... in fact, it's true whether you use c(x) and for c(x)/x above.

OpenStudy (precal):

Thanks I think I will go with AccessDenied since one of my calculus books talks about r(x) = revenue from selling x items c(x)= cost of producing the x items p(x)=r(x)-c(x) profit from selling x items if p(x)=r(x)-c(x) has a max value, it occurs at which p'(x)=0

OpenStudy (rulnick):

What I mean is, the min cost will occur at one of the boundaries in either case. Do you see it? One way the cost is convex downward, the other the cost is linear. Max profit, or min cost, will occur at x=20 or x=70.

OpenStudy (rulnick):

Well, yes, except that this is a *constrained* optimization.

OpenStudy (rulnick):

The extreme values can occur at places where the deriv is zero *or* at a boundary.

OpenStudy (accessdenied):

yeah, so you have to check those boundaries too x=20 and x=70

OpenStudy (precal):

Thanks you two have help me soo much !!!!

OpenStudy (rulnick):

welcome. good luck.